Statement of Cash Flows (Direct Method) The Wolff Companys income statement and comparative balance sheets at December 31 ofCash dividends of $37,700 were declared and paid during 2013. Plant assets were purchased for cash and bonds payable were issc. Free cash flow $ 33,800 x d. Operating-cash-flow-to-current-liabilities ratio. Round answer to two decimal places. 1.52 e.
Answer #1
Answer (a):-
Calculate the change in cash that occurred during 2013:
Changeincash=(Ending balanceofcash−Beginning balanceofcash)=$14,300−$6,500=$7,800
Change in cash during 2013 = $7,800 (increase)
Answer (b):-
Prepare the statement of cash flow on indirect method for the year ended December 31, 2013:
WOLFF COMPANY
Statement of Cash Flow
For the Year Ended December 31,2013
Particulars Amount (in $) Amount(in $)
Cash Flow from operating activities:                72,800
Net Income
Net cash provided by operating activities:
Depreciation Expenses                              22,100
Increase in accounts receivable(41600-53300)                            (11,700)
Increase in inventories(78,000-117,000)                            (39,000)
Decraese in accounts payable(9100-13000)                              (3,900)
Decraese in prepaid insurance(9100-6500)                                2,600
Decrease in income tax payable(9100-10400)                              (1,300)
incraese in wages payable(11700-7800)                                3,900              (27,300)
Net cash provided by operating activities(a)                45,500
Cash Flow from investing activities:
Purchase of Plant assets(253500-325000)                            (71,500)
Net cash used by investing activities(b)              (71,500)
Cash Flow from financing activities:
Issuance of bonds payable(169,000-97,500)                              71,500
Payment of cash dividends                            (37,700)
Net cash used by financing activities©                33,800
Net increase or (decrease) in cash (a)+(b)+©                   7,800
Cash at beginning at the year                   6,500
Cash at Closing at the year                14,300
Answer C:-
Calculate the free cash flow:
Freecashflow= (Cashflowfromoperatingactivities−Capitalexpenditure)=$45,500−$71,500= -$26,000
Therefore, the free cash flow is -$26,000.

Answer d:-

Compute the operating cash flow to current liabilities ratio

= Operating cash Flow /Average current liabities

Operating cash flow = 45,500

Average current liabilities = (accounts payable & closing +wages payable opening & closing +income tax payable opening & closing) /2

Average current liabilities = (22,100+19500+19500)/2

Average current liabilities= 30,550

operating cash flow to current liabilities ratio = 45,500/30,550

= 1.489

Answer e:-

Operating cash flow to capital expenditure = 45,500/71,500 = 0.6363

Operating cash flow to capital expenditure = operating cash flow/capital expenditure